OpenStudy (anonymous):

A sample of gas occupies a volume of 73.7 mL. As it expands, it does 133.7 J of work on its surroundings at a constant pressure of 783 torr. What is the final volume of the gas? A(S) + B(S) = C(g) A(g)+B(g) = 3C(g) A(g) + B(g) = C(g) A(s) + 2B(g) = C(g)

OpenStudy (accessdenied):

We need to understand the definition of work in terms of pressure and volume: \( \text{Work} =\Delta \left( \text{Pressure} \times \text{Volume} \right) \) We called pressure a constant, so this adjusts to (with symbols): \( W = P \Delta V \) Change in volume means the difference between Final volume and Initial volume: \( W = P \left( V_{final} - V_{initial} \right) \) We know the initial volume, the pressure and the work. So we just need to plug things in and solve for final volume from here!